How many liters of a 60% salt solution must be added to 20 liters of a 12% salt solution to produce a 40% salt solution?
Answer by mangopeeler07(462)

x=# liters
(in parentheses)=# liters
coefficient=percent

60x+12(20)=40(20+x)

Distribute
60x+240=800+40x

Subtract 240 from each side
60x=560+40x

Subtract 40x from both sides
20x=560

Divide by 20
x=28

28 liters